OpenStudy (arbyj):

someone help me solve this 12(2x - 8) = 4(5 - 2x) + 3(7 - 3x) helper will get a medal!

OpenStudy (legomyego180):

Solve for x: 12 (2 x-8)=3 (7-3 x)+4 (5-2 x) 3 (7-3 x)=21-9 x: 12 (2 x-8)=21-9 x+4 (5-2 x) 4 (5-2 x)=20-8 x: 12 (2 x-8)=-9 x+20-8 x+21 Grouping like terms, -8 x-9 x+20+21=(21+20)+(-9 x-8 x): 12 (2 x-8)=(21+20)+(-9 x-8 x) -9 x-8 x=-17 x: 12 (2 x-8)=-17 x+(21+20) 21+20=41: 12 (2 x-8)=41-17 x Expand out terms of the left hand side: 24 x-96=41-17 x Add 17 x to both sides: 24 x+17 x-96=(17 x-17 x)+41 17 x-17 x=0: 24 x+17 x-96=41 24 x+17 x=41 x: 41 x-96=41 Add 96 to both sides: 41 x+(96-96)=41+96 96-96=0: 41 x=41+96 41+96=137: 41 x=137 Divide both sides of 41 x=137 by 41: (41 x)/41=137/41 41/41=1: Answer: | | x=137/41

OpenStudy (arbyj):

thank you!!